OpenStudy (anonymous):

What is the equation of the line (in y=mx+b form) which includes points (1, 1) and (1, 5)?

jimthompson5910 (jim_thompson5910):

Are you able to graph this line?

OpenStudy (anonymous):

it doesn't say im not supposed to, but theres no graph given

jimthompson5910 (jim_thompson5910):

are you able to plot the two points on the same coordinate grid?

OpenStudy (anonymous):

same thing-- i can but im not sure if im supposed to

jimthompson5910 (jim_thompson5910):

well graphing helps see what I'm trying to point out if you graph the line, you'll see that it's a vertical line It turns out that you cannot write the equation of a vertical line in the form y = mx+b. Why? because the slope m is undefined when you have a vertical line

jimthompson5910 (jim_thompson5910):

also, if you use the slope formula to find m, you get m = (y2-y1)/(x2-x1) m = (5-1)/(1-1) m = 4/0 m = undefined

OpenStudy (anonymous):

well its multiple choice so best guess? y = 1 y = x + 1 x = 1 y = -x - 1

jimthompson5910 (jim_thompson5910):

what's the x coordinate of each point?

OpenStudy (anonymous):

1 and 1. im not dumb im just not sure of the answer

jimthompson5910 (jim_thompson5910):

correct, so x = 1 is your answer. You aren't dumb, you were probably just overthinking things.